OpenStudy (anonymous):

Determine the symmetries ( if any) of the graph of the given relation Y=7x^3 choose the correct symmetry of the graph. a. it is symmetric with respect to the origin. only_love_you_forever_268: b. It has no symmetries. c. it is symmetric with respect to the y-axis d. it is symmetric with respect to x-axis

OpenStudy (anonymous):

f(-x)=7(-x)^3=-7x^3=-f(x) this is the condition for symmetry w.r.t. the origin

OpenStudy (anonymous):

it has rotational symmetry about the origin